OpenStudy (anonymous):

Determine the zeros of the function 7x^2 + 9x = 0.

OpenStudy (anonymous):

You can take out x so that you have x(7x+9)=0 And you have x=0 and 7x+9=0 which gives you x=-9/7

OpenStudy (anonymous):

Thank you!